# From Actions to Words: Towards Abstractive-Textual Policy Summarization in RL

### v2

@@ 4.4 Generating Natural Language Summaries
3:  Let $t\leftarrow 1$
4:  $o_{t}\leftarrow$ ENVENV.RESET()
5:  $epReward\leftarrow 0$
6:  $done\leftarrow\text{False}$
7:  while $\neg done$ do
8:   $a_{t}\sim\pi(\cdot\mid o_{t})$
@@
10:   $o_{t+1},r_{t+1},done\leftarrow$ ENVENV.STEP($a_{t}$)
11:   $epReward\leftarrow epReward+r_{t+1}$
12:   Let $t\leftarrow t+1$
13:  end while
14:end for
15:formattedInputformattedInput $\leftarrow$ $P$ + FORMAT(EDED)
16:summarysummary $\leftarrow$ LLM(formattedInputformattedInput)
17:return summarysummary
5 Implementation
We showcase our method in five distinct environments from the MiniGrid Framework [49] as well as the challenging Crafter environment [50].
5.1 MiniGrid
MiniGrid is a framework for 2D based environments that feature goal-oriented tasks where agents navigate mazes, interact with objects like doors or keys, and solve challenges. We evaluate our approach in five distinct MiniGrid environments using seven different agents, each following a unique policy.
5.1.1 Captioners
To convert raw observations and actions into natural language, we designed a rule-based captioning system for the agent’s grid-based representation of the environment. This system systematically generates detailed descriptions of visible elements, emphasizing spatial relationships for navigation. Each action in the environment’s discrete action space is translated into natural language annotations, such as ‘move forward’ or ‘turn left’. We evaluated the effectiveness of this captioner by experimenting with and refining its design, ensuring the generated descriptions are accurate and comprehensive.
5.1.2 Agents’ Policy Training
We trained three qualitatively different agents for the MiniGrid-Unlock environment:
  • Goal-directed Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently. The agent can see all grid cells within a 180-degree field of vision in front of it.
  • Short-Sighted Agent: Follows a policy designed to minimize the number of steps to unlock the door efficiently but the agent’s sight is $3\times 3$.
  • Random Agent: Selects actions uniformly from the action space. We chose this policy to see how the LLM summarization process handles cases where there is no clear pattern of behavior.
Additionally, we trained four other agents in the following MiniGrid environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing. Using SySLLM, we generated a policy summary for each of these agents. All of the MiniGrid agents are trained using the PPO algorithm [51, 52]. An overview of agents’ performance is shown in Table 1. The mean reward for the first three agents in the Unlock environment shows a declining trend, which aligns with the policies that each of these agents employ. In addition, the success rate of the agents in the other environments remains consistently high. Implementation details are provided in F.
Agent Mean Reward ± SD Mean Length Success Rate
Unlock Goal-directed 0.73 ± 0.21 20.25 0.93
Unlock Short-sighted 0.41 ± 0.27 44.43 0.77
Unlock Random 0.00 ± 0.01 70.00 0.00
Dynamic Obstacles 0.78 ± 0.06 17.20 1.00
Lava Gap 0.90 ± 0.02 10.82 1.00
Red Blue Doors 0.70 ± 0.26 17.06 0.88
Crossing 0.67 ± 0.18 24.80 0.94
Table 1: Report of agents’ performance metrics for 500 episodes for the MiniGrid environments across 3 different seeds.
5.1.3 Generating Policy Summaries
For each agent, we collect an experience dataset that logs the captioned observations, actions taken by the agent, and additional relevant data. Specifically, we log 50 episodes for each agent. After collecting and formatting the data, we synthesized the narratives into coherent global policy summaries for each agent using zero-shot prompting. We employed the gpt-4-turbo model with a temperature of $0.2$ [53]. The prompt used is shown in C. In selecting the final summary, we employed a methodical approach by first generating ten distinct summaries through Line 16 of our algorithm. Each of these summaries was then embedded into a high-dimensional vector space using text-embedding-3-small, allowing us to quantify the semantic proximity of each summary to the others within this space. The centroid of these embeddings was calculated to identify the geometric mean of the summaries, representing the central tendency of the semantic dimensions explored by the model. We chose the summary that is the median (5th) closest to the centroid. This position represents a solution to the trade-off between generality and specificity. Summaries closer to the centroid tend to include more general statements that apply broadly across different scenarios, while those further away incorporate more specific details unique to particular contexts. An example of a selected summary is provided in D.
5.2 Crafter
Secondly, we evaluate SySLLM in the Crafter environment, a 2D adaptation of Minecraft. Similar to Minecraft, Crafter features a procedurally generated, partially observable world where agents can gather resources and craft artifacts structured within an achievement tree. This tree outlines all possible achievements along with their prerequisites. Crafter lacks a singular main objective, allowing agent progress to be measured by the achievements it unlocks.
5.2.1 Captioners
To convert the environment observation representations and actions in Crafter into natural-language format, we implemented a captioning system tailored to the game’s mechanics. The Observation Captioner component generates comprehensive descriptions of the player’s current state, including inventory contents, achievements, and spatial relationships to nearby resources and threats. For example, a typical observation might be transformed from raw data into ”The player is facing right. Inventory: 2 wood, 1 stone. A zombie is ahead…”. The spatial awareness system maps objects to their relative coordinates. The Action Captioner component complements this by translating action indices into meaningful descriptions like “place_table”, “move_right” or “make_wood_sword”, providing context for the agent’s decisions. Figure 3 provides an example of how the captioners translate the agent’s trajectory into natural language descriptions.
@@ 5.2.1 Captioners
Figure 3: This figure illustrates four steps from a trajectory of the Resource-Collector agent in the Crafter environment, alongside their corresponding natural language captions generated using the observation and action captioners, $C_{\text{observation}}$ and $C_{\text{action}}$. For each step, the captions describes the agent’s inventory status, the object currently in front of it, and the next action selected by the agent. A textual representation of the visible grid (highlighted in blue) is also included to reflect the agent’s local perception. Additionally, all unique achievements unlocked by the agent throughout the trajectory are summarized in red.
5.2.2 Agents’ Policy Training
We trained two qualitatively different agents for the Crafter environment. The first, named Resource-collector Agent, is a well-trained agent capable of navigating in Crafter with efficiency and skill. The agent is capable to search for food and water, defend against monsters, collect materials, and craft tools while unlocking semantically meaningful achievements along the achievement tree. This agent was trained using the DreamerV3 algorithm [54]. The second agent, a Random Agent, selects actions uniformly at random from the available action space without any learned strategy. Implementation details are provided in F.
5.2.3 Generating Policy Summaries
Similar to the MiniGrid environments, we gather an experience dataset for each agent, recording captioned observations, the actions taken by the agent, and other relevant information. Specifically, we log five episodes per agent. Due to context window limitations, each episode is summarized individually, and the resulting episode summaries are then combined into a single comprehensive summary. The prompts used are shown in E.
5.3 Insights From Summaries
Using the SySLLM method, we produced summaries that highlight essential aspects of each agent’s decision-making and behavior. This section outlines key insights derived from these summaries, offering a clear depiction of the agents’ policies and significant action patterns. Table 3 presents an overview of these insights for the MiniGrid agents, while Table 4 provides an overview for the Crafter agents. These insights showcase the alignment of SySLLM-generated summaries with task objectives and illustrate how agents perform across diverse environments. For the goal-directed agent in the Unlock environment, Insight (2) states that the agent “demonstrates a pattern of turning towards the nearest key or door once identified, suggesting a straightforward heuristic of minimizing distance to the target.” This reflects the agent’s efficient and consistent behavior, dynamically adjusting its path to achieve objectives with minimal steps. Insight (3) quantifies this efficiency, noting that episodes are completed within 15-25 steps on average, often achieving near-maximum rewards. These observations are validated by performance metrics in Table 1, which report a mean episode length of 20.25 and a high mean reward of 0.71. Together, these results highlight SySLLM’s ability to capture both qualitative patterns and quantitative performance metrics.
In the Dynamic Obstacles environment, SySLLM captures the agent’s ability to navigate challenges effectively. Insight (1) states that the agent “consistently demonstrates the ability to avoid obstacles (blue balls) by making turns or moving forward when the path is clear.” This highlights its proficiency in avoiding collisions, a critical success factor. However, Insight (3) notes inconsistent behavior when navigating closely spaced obstacles, with the agent occasionally taking longer routes or making unnecessary turns. These observations demonstrate SySLLM’s ability to provide a nuanced understanding of both the agent’s strengths in straightforward scenarios and its struggles in more complex configurations. For the Resource-collector agent in the Crafter environment, SySLLM effectively captures the agent’s behavior as a resource-collector, emphasizing its focus on gathering essential materials. Insight (1) highlights that “The agent consistently prioritizes acquiring fundamental resources such as wood, stone, and drink, which are critical for crafting tools and sustaining basic survival metrics.” Furthermore, SySLLM captures the agent’s achievement-oriented tendencies while identifying limitations in unlocking advanced achievements. As stated in Insight (4), “The agent reliably achieves milestones related to resource collection and basic tool crafting but demonstrates difficulty in attaining more advanced achievements, suggesting potential areas for improvement.” These observations are validated by performance metrics presented in Table 2, which report a mean reward of 10.42—indicating relatively high performance—and a mean of 11.33 unique achievements unlocked, reinforcing the agent’s proficiency in foundational tasks while highlighting room for growth in advanced capabilities.
Agent Mean Reward ± SD Mean Length Mean Unique Achievements
Resource-Collector 10.43 ± 2.11 234.6 11.33
Random 1.39 ± 1.19 164.4 2.29
Table 2: Report of agents’ performance metrics for 100 episodes for the Crafter environment across 3 different seeds.
Agent Insights from Summaries
Unlock
Goal–
directed
@@ 5.3 Insights From Summaries
(1) The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
@@
(2) It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
@@
(3) The agent completes episodes efficiently, averaging 15–25 steps with near-maximal cumulative rewards.
Unlock
Short–
sighted
@@
(1) The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
@@
(2) It identifies keys and doors efficiently, maneuvering toward and using them correctly.
@@
(3) Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
Unlock
Random
@@
(1) The agent exhibits unstructured behavior, often repeating unnecessary actions.
@@
(2) It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
@@
Lava Gap (1) The agent consistently avoids lava, demonstrating awareness of environmental hazards.
@@
(2) Upon encountering an obstacle, it either turns or moves in the opposite direction.
@@
(1) The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
@@
(2) It successfully interacts with doors in a structured sequence, adhering to task constraints.
Crossing (1) The agent moves towards the green goal once it enters its field of vision, adjusting its path accordingly.
(2) It avoids collisions with walls through timely directional changes.
Dynamic
Obstacles
(1) The agent effectively avoids moving obstacles (blue balls) by adjusting its movement.
(2) It identifies objects in its field of vision and makes informed navigation decisions.
(3) In dense obstacle scenarios, occasional inefficiencies or unnecessary turns are observed.
Table 3: Insights from agents’ SySLLM summaries in the MiniGrid environments.
Agent Insights from Summaries
Crafter
Resource-collector
(1) Strong Focus on Resource Collection: The agent consistently prioritizes gathering essential resources such as wood, stone, and drink, which are foundational for crafting tools and maintaining basic survival metrics.
(2) Effective Basic Tool Crafting: Regular crafting of basic tools like wood pickaxes and swords enables the agent to enhance resource collection and engage in occasional combat.
(3) Achievement Unlocking: The agent reliably unlocks achievements related to resource collection and basic tool crafting but struggles with more advanced achievements, highlighting a potential area for improvement.
(4) Moderate Combat Engagement: The agent occasionally engages with zombies, using crafted tools for defense, showing moderate adaptability to threats but limited combat readiness overall.
(5) Predictable Behavior: Episodes are characterized by high consistency in resource collection and basic tool crafting actions, with low variance across episodes.
(6) Inconsistent Health and Exploration Management: While the agent effectively manages drink levels, it shows less consistency in food and health management and sacrifices exploration efficiency for achievement unlocking.
Crafter
Random
(1) Ineffective Crafting and Resource Management: The agent frequently attempts crafting without the necessary resources or understanding of prerequisites, leading to repeated failures and minimal progress in achieving complex objectives.
(2) Poor Survival Strategy: The agent consistently demonstrates ineffective survival behavior, including health depletion and poor management of food and drink levels, which hampers its ability to sustain itself in the game environment.
(3) Limited Achievement Progression: While the agent reliably unlocks basic achievements like ‘wake_up’ and ‘collect_sapling’, it struggles to achieve more complex milestones that require crafting, resource management, or combat engagement.
(4) Repetitive and Ineffective Actions: Episodes are marked by high frequencies of ‘noop’ actions and repetitive failed attempts at crafting, reflecting a lack of strategic adaptation and learning from past failures.
(5) Lack of Combat Engagement: Lack of Combat Engagement: The agent shows minimal engagement with combat mechanics and fails to defend effectively against threats such as zombies and skeletons.
(6) Predictable Behavior: Across episodes, the agent exhibits consistent, ineffective patterns of action, suggesting significant limitations in its decision-making processes and adaptability.
Table 4: Insights from agents’ SySLLM summaries in the Crafter environment.
6 Expert Evaluation of Summaries
To evaluate the quality and alignment of SySLLM summaries, we conducted evaluations with RL experts. Specifically, six graduate students actively engaged in RL research with experience in training agents were recruited. These experts were divided into two groups: Experts 1, 2, and 3 evaluated the MiniGrid-Unlock goal-directed and short-sighted agents, while Experts 4, 5, and 6 evaluated the the rest of the MiniGrid agents. Experts 1,2, and 6 evaluated the agents from the Crafter environment. Each expert was shown videos approximately 120 seconds long, depicting the agents’ behaviors in the environment. Based on these videos, the experts were tasked with summarizing the agents’ behaviors. To ensure a fair comparison with SySLLM-generated summaries, we provided the experts with instructions that aligned closely with the prompts used for the language model (see G for detailed instructions).
To evaluate how much of the experts’ observations were captured by the SySLLM summaries, we used a recall metric. This metric quantified the extent to which the SySLLM summary covered the points identified by the experts. Each expert summary and the SySLLM summary were broken down into lists of key points, and a scoring system was applied: 1 point for a match with a SySLLM key point, 0.5 points for a partial match, and 0 points for no match. The recall scores, as shown in Table 5, ranges from 0.687 for the Unlock goal-directed agent, indicating moderate agreement, to 0.914 for the Crossing agent, reflecting a near-complete capture of expert insights. The overall mean recall score across all agents is 0.840, demonstrating substantial coverage of the experts’ points by the SySLLM summaries.
In the second evaluation phase, the experts were presented with key points from the SySLLM summaries that they had not mentioned in their own summaries. They were tasked with classifying these points as Matched, Partially Matched, or Not Matched based on their agreement with the content. Using the same scoring system as for recall (Matched = 1, Partially Matched = 0.5, Not Matched = 0), we calculated a precision score for each summary. The precision score measures how many of the SySLLM summary points were validated as correct by the experts. Importantly, the precision calculation includes all SySLLM key points, encompassing both the points classified in this phase and those already evaluated during the recall assessment. The precision scores, as shown in Table 5, range from 0.769 (Dynamic Obstacles agent) to 0.864 (Short-Sighted agent), with an overall mean precision score of 0.839 across all agents. In addition to quantifying correctness, this metric offers a way to evaluate potential hallucinations by identifying points that may not align with expert observations. The high precision scores suggest that hallucinations were minimal in our evaluations.
While instances of hallucinated content in SySLLM summaries were rare, they underscore the inherent challenges in summarizing RL agent behaviors using LLMs. One example from the Crossing agent summary: “The agent frequently checks for walls in its path and adjacent tiles.”. This point is classified as “not matched” by all experts, indicating a possible hallucination.
To assess the reliability of expert annotations, we report both percentage agreement and Gwet’s AC1–metrics appropriate for imbalanced labeling tasks. The mean agreement reached 70%, and Gwet’s AC1 indicated substantial reliability (mean = 0.72), reflecting strong consistency across annotators. These results not only validate the quality of the annotations, but also provide important context for interpreting model performance: if human experts do not fully agree, perfect alignment between model and human labels is neither expected nor realistic. In this light, the model’s ability to approximate expert consensus is particularly notable.
Taken together, the integration of recall and precision metrics demonstrates SySLLM’s capability to produce comprehensive and accurate summaries that closely align with expert observations.
Agent Expert Recall Precision Mean Recall Mean Precision
Unlock
Goal-Directed
1 0.500 0.864 0.687 0.864
2 0.643 0.864
3 0.917 0.864
Unlock
Short-Sighted
1 0.800 0.846 0.878 0.839
2 0.833 0.807
3 1.000 0.923
Dynamic
Obstacles
4 0.583 0.692 0.739 0.769
5 0.833 0.692
6 0.800 0.923
Lava Gap
4 0.667 0.769 0.794 0.811
5 0.786 0.846
6 0.929 0.818
Red Blue
Doors
4 0.857 0.767 0.871 0.834
5 0.857 0.867
6 0.900 0.867

### v3

@@ 4.3. Abstractive Summarization
3:  $t\leftarrow 1$; $o_{t}\leftarrow ENV.\texttt{RESET}()$; $epReward\leftarrow 0$; $done\leftarrow\textbf{False}$
4:  while $\neg done$ do
5:   $a_{t}\sim\pi(\cdot\mid o_{t})$
6:   $TEB.\texttt{ADD}\big(\,C_{\text{obs}}(o_{t}),\,C_{\text{act}}(a_{t}),\,epReward,\,i\,\big)$
7:   $(o_{t+1},r_{t+1},done)\leftarrow ENV.\texttt{STEP}(a_{t})$
8:   $epReward\leftarrow epReward+r_{t+1}$; $t\leftarrow t+1$
9:   $o_{t}\leftarrow o_{t+1}$
10:  end while
11:end for
@@
13:return $T$
Algorithm 2 HierarchicalSummarize
Input: textual experience buffer subset $D$, token budget $\kappa$, prompt $P$, #candidates $K$, embedding model $\phi$
Output: summary $T\in\mathcal{T}$
1:if $\texttt{TOKENS}(\texttt{FORMAT}(D))\leq\kappa$ then
2:  $X\leftarrow P\;+\;\texttt{FORMAT}(D)$
3:  $\{T_{1},\ldots,T_{K}\}\leftarrow\texttt{LLM.SAMPLE}(X,K)$
4:  $E_{i}\leftarrow\phi(T_{i})\in\mathbb{R}^{d}\quad\forall i\in\{1,\ldots,K\}$
5:  $c\leftarrow\frac{1}{K}\sum_{i=1}^{K}E_{i}$
6:  $d_{i}\leftarrow\lVert E_{i}-c\rVert_{2}\quad\forall i$;   $j^{\star}\leftarrow\arg\min_{j}\left|\,\texttt{rank}(d_{j})-\frac{K}{2}\,\right|$
7:  return $T_{j^{\star}}$
8:else
9:  $\{D^{(1)},\ldots,D^{(M)}\}\leftarrow\texttt{PARTITION}(D,\kappa)$ $\triangleright$ disjoint, each fits
10:  for $m=1$ to $M$ do
11:   $S_{m}\leftarrow\textsc{HierarchicalSummarize}(D^{(m)},\kappa,P,K,\phi)$
12:  end for
[…]
14:end if
5. Experimental Setup
We evaluate SySLLM across controlled reinforcement learning environments that capture both simple and complex agent behaviors. Specifically, we apply our framework to five environments from the MiniGrid suite chevalier2021minimalistic and the more challenging Crafter environment hafner2021crafter. These environments were selected to span a range of task structures, observation modalities, and policy complexities. We use the gpt-4-turbo model with a temperature of 0.5 achiam2023gpt, and the text-embedding-3-small model openai2024embedding as the embedding function.
5.1. MiniGrid
MiniGrid is a grid-world framework where agents perform goal-directed navigation and object-interaction tasks under partial observability. We instantiate seven agents across five environments, ensuring diversity in both policy performance and behavioral style.
Captioners. To construct the Textual Experience Buffer (TEB), we implement a rule-based captioning system that maps raw grid observations and discrete actions into structured natural-language descriptions. Observation Captioner ($C_{\text{obs}}$): generates textual descriptions of visible elements, including object types (e.g., keys, doors, obstacles) and their spatial relationships relative to the agent. Action Captioner ($C_{\text{act}}$): translates the discrete action set (e.g., turn_left, move_forward) into natural-language strings. Agents and Training We trained three agents with qualitatively distinct policies in the MiniGrid-Unlock environment:
  • Goal-directed agent: wide $7\times 7$ observation window, optimized to minimize the steps to unlock the door.
  • Short-sighted agent: restricted $3\times 3$ observation window, leading to more myopic strategies.
  • Random agent: selects actions uniformly at random, providing a non-structured behavioral baseline.
In addition, we trained one agent each in four further environments: Dynamic Obstacles, Lava Gap, Red-Blue Doors, and Crossing. All MiniGrid agents were trained using PPO schulman2017proximal; stable-baselines3 for 1M timesteps per seed for three random seeds. Agent performance statistics are reported in Table 1. The complete hyperparameters are provided in the Appendix F.
TEB Collection and Summarization. For each agent, we collect $50$ evaluation episodes to construct the TEB. Each buffer is formatted into the structured prompt and processed as described in Section 4.3, where candidate summaries are generated via stochastic decoding and post-processed for abstractive summarization. Prompt templates used for the MiniGrid suite domain are listed in Appendix C.
5.2. Crafter
Crafter is a 2D, partially observable world inspired by Minecraft, featuring procedurally generated maps, resource gathering, crafting, and an achievement tree that defines agent progress.
Captioners To adapt the captioning system to Crafter, we extend the MiniGrid captioner to encode: Observation Captioner: inventory contents, spatial relations to nearby resources and threats, current health and stamina, and unlocked achievements. Action Captioner: maps the Crafter action set into textual forms, e.g., “move_right”, “place_table”, “make_wood_sword”. Figure 3 shows an example trajectory and its textualized representation.
Agents and Training We train two agents with distinct behavior:
  • Resource-Collector agent: trained with DreamerV3 hafner2023dreamerv3, capable of sustained survival, resource collection, and the crafting of basic tools.
  • Random agent: uniformly samples from the action set, serving as a baseline with no structured policy.
TEB Collection and Summarization. For each agent, we log $5$ evaluation episodes. Due to the long horizon of Crafter, the TEB for each episode can exceed the LLM’s context length. In this case, we apply the hierarchical summarization strategy described in Section 4.3. Prompt templates used for the Crafter domain are listed in Appendix E.
[…]
Crafter captioning
@@ 5.2. Crafter
Figure 3. Four steps from a trajectory of the Resource-Collector agent in the Crafter environment, alongside their corresponding captions generated using the observation and action captioners. For each step, the captions describe the agent’s inventory status, the object currently in front of it, and the next action selected by the agent. A textual representation of the visible grid (highlighted in blue) is also included to reflect the agent’s local perception. Additionally, all unique achievements unlocked by the agent throughout the trajectory are summarized in red.
Env. Agent Mean Reward ± SD Mean Length Success / Achievements
MiniGrid Unlock Goal-directed 0.73 ± 0.21 20.25 Success 0.93
Unlock Short-sighted 0.41 ± 0.27 44.43 Success 0.77
Unlock Random 0.00 ± 0.01 70.00 Success 0.00
Dynamic Obstacles 0.78 ± 0.06 17.20 Success 1.00
Lava Gap 0.90 ± 0.02 10.82 Success 1.00
Red Blue Doors 0.70 ± 0.26 17.06 Success 0.88
Crossing 0.67 ± 0.18 24.80 Success 0.94
Crafter Resource-Collector 10.43 ± 2.11 234.6 11.33 achievements
Random 1.39 ± 1.19 164.4 2.29 achievements
Table 1. Performance metrics across MiniGrid (500 episodes, 3 seeds) and Crafter (100 episodes, 3 seeds). For MiniGrid, performance is measured by success rate; for Crafter, by the number of unique achievements unlocked.
6. Illustrative Policy Summaries
To illustrate the summaries produced by SySLLM, we present structured case studies that connect textual motifs to quantitative evidence and then expand them into detailed narrative accounts. This combination shows that the framework produces explanations that are not only linguistically coherent but also grounded in the actual behavior of the agents. A full summary of the Unlock Goal-Directed agent is included in Appendix D.
Structured Case Studies. Table 2 reports representative agents in MiniGrid and Crafter. For each, we show: (i) the central claim extracted by SySLLM, and (ii) the quantitative metrics that support or refute this claim. The alignment between narrative motifs and behavioral statistics illustrates that SySLLM captures recurring strategies and limitations in a manner consistent with ground-truth agent performance. For example, the Unlock Goal-Directed agent’s motif of “turning toward the nearest key or door” is supported by a $0.93$ success rate and shorter episode lengths, while the Random agent’s “lack of coherent strategy” corresponds to maximal episode lengths and zero success rate. In Crafter, the Resource-Collector agent’s emphasis on survival resources is matched by high reward and achievement counts, whereas the Random agent’s incoherence is validated by minimal achievement progression.
Env. Agent SySLLM Summary Claim Quantitative Alignment
MiniGrid Unlock Goal-directed Turns toward nearest key/door Success $0.93$, mean length $20.25$
Unlock Random No clear strategy Success $0.00$, max length episodes
Dynamic Obstacles Avoids obstacles, occasional inefficiency Success $1.00$, higher variance in length
Lava Gap Jumps gap consistently Success $1.00$, mean length $10.82$
Crafter Resource-Collector Focused, survival-oriented strategy; Unlocks early achievements Reward $10.43$, $11.33$ unique achievements
Random Sporadic, incoherent progression Reward $1.39$, $2.29$ unique achievements
Table 2. Structured case studies of SySLLM summaries across environments. Extracted motifs are validated against quantitative metrics, showing alignment between narrative claims and behavioral statistics.
Expanded Narrative Insights. Beyond single-sentence claims, SySLLM produces multi-faceted descriptions of agent behavior. Figure 4 presents detailed motifs across environments, highlighting how the framework abstracts local decisions into global patterns. These narratives capture not only the strengths of agents (e.g. consistent lava avoidance, structured door-order strategies) but also nuanced inefficiencies (e.g., unnecessary turns in dense obstacle scenarios, repetitive failed crafting attempts in Crafter). Such fine-grained insights demonstrate SySLLM’s ability to diagnose effective heuristics and characteristic failure modes.
Unlock – Goal-directed
  1. (1)
@@ 6. Illustrative Policy Summaries
    The agent effectively identifies keys and adjusts its path based on their relative position, shifting focus to unlocking the door.
  2. (2)
@@
    It consistently turns towards the nearest key or door, minimizing distance, which remains consistent across episodes.
  3. (3)
@@
    The agent completes episodes efficiently, averaging 15–25 steps with near-maximal cumulative rewards.
Unlock – Short-sighted
  1. (1)
@@
    The agent follows a right-wall method, moving forward until encountering an obstacle before turning.
  2. (2)
@@
    It identifies keys and doors efficiently, maneuvering toward and using them correctly.
  3. (3)
@@
    Decisions are heavily influenced by its immediate field of vision, reacting only to nearby objects.
Unlock – Random
  1. (1)
@@
    The agent exhibits unstructured behavior, often repeating unnecessary actions.
  2. (2)
@@
    It frequently toggles doors multiple times or picks up and drops keys without using them effectively.
Lava Gap
  1. (1)
@@
    The agent consistently avoids lava, demonstrating awareness of environmental hazards.
  2. (2)
@@
    Upon encountering an obstacle, it either turns or moves in the opposite direction.
@@
    The agent prioritizes opening the red door before the blue door, optimizing reward accumulation.
  2. (2)
@@
    It successfully interacts with doors in a structured sequence, adhering to task constraints.